what is proportationugyuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuutydrudsuftttttttttttttttigbyunnnnnnnnnnnnnnnnnnnnnnnnnnnnnnnnnnnnbbbbbbbbbbbbbbbbbu8ityvu7rrrrrrcccccccccccccccccc77777777

Grade:6

1 Answers

Harshit Singh
3 years ago
Dear Student

Proportion says that two ratios (or fractions) are equal.
It can be written in two ways:
1. as two equal fractions a/b = c/d;
2.using a colon, a:b = c:d.

ex – 2/4=1/2

Thanks